The short-run cost function of a company = TC=200+55y( tc,y measured in thousands)

a) what is the fixed cost ?

b) ıf the company produces 100.000 units of output, what would be the AVC ?

c) what would be its marginal cost (y=100.000)

d) what would be its average fixed cost (y=100.000)

e)suppose that company expounds, its production capacity, therefore, fixed cost increases by $50.000, but its variable cost falls to $ 45.000 per1000 units write the new cost equation?

f)Suppose the company borrows money and expands its factory. Its fixed cost rises by $50.000, but its variable cost falls to $45.000 per 1000 units. The cost of interest ( ) also enters into the equation. Each 1 -point increase in the interest rate raises costs by $3000 Write the new cost equation


1
Expert's answer
2022-03-25T15:07:58-0400

a) what is the fixed cost ?

"FC = 200" (in thousands)


b) ıf the company produces 100.000 units of output, what would be the AVC ?

The variable cost here is "55y"

"AVC =\frac{\Delta VC}{Y} =\frac{ 55y}{100}=0.55y"


c) what would be its marginal cost (y=100.000)

TC=200+55y

"MC = \frac{\Delta TC}{Y} = 55" (in thousands)


d) what would be its average fixed cost (y=100.000)

"AFC = \frac{\Delta FC}{Y} = \frac{200000}{100}= 2000"


e)suppose that company expounds, its production capacity, therefore, fixed cost increases by $50.000, but its variable cost falls to $ 45.000 per1000 units write the new cost equation?

"TC = 250+45y"